Walt made an extra $5000 last year from a part-time job. He invested part of the money at 6% and the rest at 10%. He made a total of $420 in interest. How much was invested at 10%?
Answer by checkley77(12844)

.10x+.06(5,000-x)=420
.10x+300-.06x=420
.04x=420-300
.04x=120
x=120/.04
x=3,000 amount invested @ 10%.
Proof:
.10*3,000+.06(5,000-3,000)=420
300+.06*2,000=420
300+120=420
420=420